OpenStudy (anonymous):

√(x+2)+4=X show the work please :)

OpenStudy (saifoo.khan):

\[\sqrt{x+2} = 4-x\] \[x+2 = (4-x)^2\] solving this will give u two values, 7 and 2. now insert them back to check. Final answer will be 7.

OpenStudy (saifoo.khan):

just 7.